# ShortcutsBench: A Large-Scale Real-world Benchmark for API-based Agents

### v1

@@ 1 Introduction
Some existing benchmarks/datasets (Huang et al., 2023; Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024; Hao et al., 2024) have attempted to evaluate API-based agents. However, they have three limitations: First, the APIs (a.k. tools available to the agent) lack richness, and the queries (a.k. the task to the agent) lack complexity (Table 1). They either involve a limited number of APIs (Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024) or cover only a few numbers of apps (Patil et al., 2023; Tang et al., 2023; Schick et al., 2024; Hao et al., 2024) (an app may have one or multiple APIs), or the difficulties of the queries are limited in a narrow range, with the average action length ranges from $1$ to $5.9$ (Table  1). This lack of richness and complexity makes it difficult to effectively distinguish the capabilities of different agents, particularly on larger and smarter LLMs like Gemini-1.5 Pro DeepMind (2024) and GPT-3.5-turbo OpenAI (2023a). Second, the APIs lack realism as they may be manually crafted, and the queries fail to reflect actual user demands since they mey be either created by hand or generated directly by ChatGPT OpenAI (2023a) without verifying real user demands (Table 1). Moreover, they only cover the evaluation of API selection, lacking a study on API parameter filling (Table 1). Efficient and accurate parameter filling is essential for an agent to complete tasks successfully. Thirdly, they don’t adequately evaluate the agent’s ability to ask systems or the users for the necessary information to resolve the missing information for solving the queries (Table 1). This is crucial as a user’s query may be implicit or may not provide all the information an agent needs to solve the task (Qin et al., 2023; Qian et al., 2024) effectively.
@@
# Queries 7628 21127 12657 17002 3938 274 2726 1530
# Avg APIs 7.86 1.0 2.3 1.0 1.0 2.1 5.4 13.0
# Avg Actions 21.46 1.0 4.0 1.0 1.0 2.2 5.9 1.0
@@ 3.1 Dataset Acquisition
Figure 1 shows the data acquisition process. We first use search engines to identify popular public shortcut-sharing sites ①. We totally find $14$ sites such as Shortcuts Gallery (Gallery, 2024) and SSPai (SSPai, 2024). Then we crawled these sites to obtain fields such as “shortcut name”, “function description”, “shortcut type”, and “iCloud link” ②. After deduplicating based on “iCloud link” (Apple, 2024c), we got the source files of all $8675$ shortcuts (Apple, 2024b) ③. Subsequently, we extracted “app name” using the  field WFWorkflowActionIdentifier in the source file like com.apple.Music, and then downloaded related apps ④ from various sources: (1) third-party apps from the “macOS App Store” or the official website of the app, (2) non-system apps (uninstallable) like Keynote from path /Applications/ on macOS, (3) system apps like Reminders (uninstallable) from path /System/Application/ on macOS, (4) iOS apps from the “iOS App Store”, (5) Shortcuts app itself from path /System/Library/PrivateFrameworks/WorkflowKit.framework/ on macOS. During the downloading, we also excluded some legacy apps and $12$ paid apps. For more details about the whole acquisition process, please refer to Appendix.
Then we managed to extract APIs from the downloaded apps ⑤. The APIs are mainly from intent definition file \${filename}.actionsdata from AppIntent (Inc., 2024b; 2022) framework, \${filename}.intentdefinition from SiriKit (Inc., 2024c; 2022) framework, and WFActions.json from system path /System/Library/PrivateFrameworks/WorkflowKit.framework/ on macOS. We extracted all APIs involved in the app’s shortcuts. During the extraction, due to the existence of multi-platform versions of most apps like Day One (Day One, 2024a; b), and the presence of similar and duplicate API definition files from the same and different frameworks (Inc., 2024c; b; 2022), we perform deduplication based on manually crafted rules. For more details about deduplication, please refer to Appendix. Finally, as shown in Table 1, we get $88$ apps from various categories such as “Health & Fitness”(iTunes App Store, 2024b), “Developer Tools” (iTunes App Store, 2024a), and “Lifestyle” (iTunes App Store, 2024c). These apps in total include $1414$ APIs, including all of $556$ APIs involved in $7628$ shortcuts.
@@ 3.2 Dataset Construction
(“a” in Figure 2) include the “API description” (“a.4”), “API name” (“a.1”), “parameter names” (“a.2”), “parameter types” (“a.1”), “default value” (“a.2”), “return value type” (“a.3”), and “return value name” (“a.3”). The field names in [...] in Figure 2 correspond to the field names in \${filename}.actionsdata, \${filename}.intentdefinition, and WFActions.json. For details, please refer to the Appendix. In existing benchmarks/datasets (Table 1), the “parameter types” (“a.1”) and “return value types” (“a.3”) are composed of primitive data types such as int and string. In addition to primitive data types, APIs in ShortcutsBench also include “enum” or “advanced data types”. Enum is composed of “the class name” and “the possible value”, with each value equipping a “value name”. We also provide the agent with a description of the “enum” in the API information. Advanced data types, such as the model (“a.1”) in app chatGPT (OpenAI, 2024), include three String types named identifier, title, and subtitle. We can comprehend them through their “type name” and “type description”.
@@
Generating queries. As shown in Figure 1, existing works construct query and action sequences based on available APIs. In contrast, we construct queries based on existing action sequences and APIs. When constructing a query for a specific action sequence, we need to understand the functional description of the action sequence (“e” in Figure 2) and detailed information about the involved APIs (“a” in Figure 2). With this information, we can generate higher-quality queries. To ensure the quality of the generated queries, we also leverage the unique advantage of shortcuts: the natural language workflow descriptions (“b.7” / “c” in Figure 2). By inputting these intuitive natural language descriptions into an LLM, we can generate more accurate queries. When generating queries, we also require the model to naturally include primitive data type parameters and enum data type needed for API calls in generated queries. This helps us evaluate the agent’s ability to fill in primitive parameters in Section 3.3.2. To ensure the quality of generated queries, we use the state-of-the-art LLM, GPT-4o (OpenAI, 2024b), to generate the queries. The prompt templates we used to generate queries can be found in the Appendix.
@@ 3.3 Task Definition and Metrics
Further Processing. Considering the context limitations of LLMs, we excluded shortcuts longer than $30$ and parts using the API is.workflow.actions.runworkflow to call other shortcuts. While these shortcuts remain in our open-source dataset, they will not be included in the subsequent evaluation. We aim to study the performance of agents on queries of varying difficulties. As shown in Table 2, we categorize ShortcutsBench into $4$ difficulty levels and $8$ task types based on $\left|aseq_{i}\right|$ and “shortcut type” (Section 3.1), respectively. For more details, please refer to the Appendix. When calculating the length, for branching actions like is.workflow.actions.conditional, we consider the longest branch as the length. Additionally, we ignore the lengths of looping actions like is.workflow.actions.repeat.count and special actions such as is.workflow.actions.comment. Due to the presence of branching actions, the average number of APIs involved when $p=1$ is greater than one, specifically $1.76$. For a detailed process, please refer to the Appendix. The number of shortcuts in each level is denoted as $n_{p}$. Each query and action sequence is referred to as $q_{p,i}$ and $aseq_{p,i}$, with $1\leq p\leq 4$ and $1\leq i\leq n_{p}$.
@@ 3.3.1 Performance about API Selection
Following existing work (Huang et al., 2023; Patil et al., 2023; Li et al., 2023b; Xu et al., 2023; Schick et al., 2024; Hao et al., 2024), we use the accuracy of API selection as the metric. The accuracy is calculated as the number of correct API selections $m_{p}$ divided by $n_{p}$. Specifically, each time we predict an action $b_{j},1\leq j\leq|aseq_{i}|$, we provide the agent with all the correct historical actions $\{a_{1},a_{2},...,a_{j-1}\}$. We then require the agent to predict the next action. All actions predicted by the agent form the prediction sequence $bseq_{p,i}$. This method is similar to the next token prediction (NTP) in LLMs, effectively preventing a cascade of errors in subsequent action predictions due to a single incorrect prediction. During the prediction, when encountering special actions such as branching and looping, we skip predicting these actions and directly add them to the historical actions. For more details, please refer to Appendix. We chose API selection accuracy over the final result for the following two additional reasons:
@@ 3.3.2 Effectiveness of API Parameter Value Filling
  • Static Parameters Preset: These are static parameters that users provide as default inputs of the action. These static parameters typically include primitive data types such as String and Integer, as well as custom Enum defined by app developers. When the query explicitly specifies a parameter that can be used as a static parameter, we expect the agent to accurately fill in the parameter values according to the user’s query and the API’s definition. When generating queries, we have already required the LLM to naturally include primitive and enumerated data types (Section 3.2). To further ensure that the corresponding parameters are indeed included in the queries during evaluation, we used the LLM to filter these parameters further, ensuring their presence in the queries. Detailed prompts can be found in the Appendix.
@@ 3.3.3 Recognition of Need for Input
In this section, we aim to investigate the ability of existing API-based agents to ask systems or users for necessary information to resolve the missing information. This missing information can come from the system like clipboard (Clipboard), input files (ExtensionInput), and the current date (CurrentDate) or from the user (Ask(Inc., 2024a). For example, a parameter named tags is usually represented in a shortcut as "tags":{"Value":{"Type": "Ask"}}, where "Type": "Ask" indicates that the parameter will prompt the user for input. For a detailed process, please refer to Appendix. We use the proportion of correctly identified parameters to evaluate the agent’s ability to recognize the need for input from the system or the user. Let $n_{s}$ be the number of queries, $aska_{i}$ be the number of times the need from the system or the user appears in $aseq_{i}$, $askt_{i}$ be the number of times the need from the system or the user appears in $bseq_{i}$. The accuracy of ask for necessary information can be calculated as $Acc_{afni}=askt_{i}/aska_{i}$.
@@ 4.1 Setup
Prompt Template. Following existing work (Huang et al., 2023; Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024; Hao et al., 2024), we slightly modified the ReACT (Yao et al., 2022) templates to construct the API-based agents. Specifically, we added prompts related to shortcuts, such as the types of fillable parameters and the meanings of special statements like branches. For all $3$ research questions (RQs), we use the same prompt templates. An agent should correctly select APIs, fill in parameters, and be aware of the need to request necessary information from the system or user at appropriate times. Please refer to Appendix for more details.
@@ 4.2 Result Analysis
  • Agents built using open-source LLMs now perform comparably to closed-source models on lower-difficulty tasks but still lag on higher-difficulty tasks. From Figure 4 we know that open-source LLMs $>=70$B match the performance of closed-source LLMs from the first $3$ difficulty tasks, significantly outperforming GPT-3.5-turbo. However, they still lag behind closed-source LLMs in handling complex tasks at the $4$-th level. Moreover, the price of open-source LLMs is significantly lower than that of GPT-3.5-turbo. For more details, please refer to Appendix.

### v2

@@ 1 Introduction
Some existing benchmarks / datasets (Huang et al., 2023; Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024; Hao et al., 2024) have attempted to evaluate API-based agents. However, they have three limitations: First, the APIs (a.k. tools available to the agent) lack richness, and the queries (a.k. the task to the agent) lack complexity (Table 1). They either involve a limited number of APIs (Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024) or cover only a few numbers of apps (Patil et al., 2023; Tang et al., 2023; Schick et al., 2024; Hao et al., 2024) (an app may have one or multiple APIs), or the difficulties of the queries are limited in a narrow range, with the average action length ranges from $1$ to $5.9$ (Table  1). This lack of richness and complexity makes it difficult to effectively distinguish the capabilities of different agents, particularly on larger and smarter LLMs like Gemini-1.5 Pro DeepMind (2024) and GPT-3.5-turbo OpenAI (2023a). Second, the APIs lack realism as they may be manually crafted, and the queries fail to reflect actual user demands since they may be either created by hand or generated directly by ChatGPT OpenAI (2023a) without verifying real user demands (Table 1). Moreover, they only cover the evaluation of API selection, lacking a study on API parameter filling (Table 1). Efficient and accurate parameter filling is essential for an agent to complete tasks successfully. Thirdly, they don’t adequately evaluate the agent’s ability to ask systems or the users for the necessary information to resolve the missing information for solving the queries (Table 1). This is crucial as a user’s query may be implicit or may not provide all the information an agent needs to solve the task (Qin et al., 2023; Qian et al., 2024) effectively.
@@
# Queries 7627 21127 12657 17002 3938 274 2726 1530
# Avg APIs 9.62 1.0 2.3 1.0 1.0 2.1 5.4 13.0
# Avg Actions 21.62 1.0 4.0 1.0 1.0 2.2 5.9 1.0
@@ 3.1 Dataset Acquisition
Figure 1 shows the data acquisition process. We first use search engines to identify popular public shortcut-sharing sites ①. We totally find $14$ sites such as Shortcuts Gallery (Gallery, 2024) and SSPai (SSPai, 2024). Then we crawled these sites to obtain fields such as “shortcut name”, “function description”, “shortcut type”, and “iCloud link” ②. After deduplicating based on “iCloud link” (Apple, 2024c), we got the source files of all $8675$ shortcuts (Apple, 2024b) ③. Subsequently, we extracted “app name” using the  field WFWorkflowActionIdentifier in the source file like com.openai.chat.AskIntent, and then downloaded related apps ④ from various sources: (1) third-party apps from the “macOS App Store” or the official website of the app, (2) non-system apps (uninstallable) like Keynote from path /Applications/ on macOS, (3) system apps like Reminders (uninstallable) from path /System/Application/ on macOS, (4) iOS apps from the “iOS App Store”, (5) Shortcuts app itself from path /System/Library/PrivateFrameworks/WorkflowKit.framework/ on macOS. During the downloading, we also excluded some legacy apps and $12$ paid apps. For more details about the whole acquisition process, please refer to Appendix A.1.
Then we managed to extract APIs from the downloaded apps ⑤. The APIs are mainly from intent definition file \${filename}.actionsdata from AppIntent (Apple-Inc., 2024b; 2022; 2023) framework, \${filename}.intentdefinition from SiriKit (Apple-Inc., 2024e; 2022; 2023; c) framework, and WFActions.json from system path /System/Library/PrivateFrameworks/WorkflowKit.framework/ on macOS. We extracted all APIs involved in the app’s shortcuts. During the extraction, we perform deduplication of APIs based on manually crafted rules. Finally, as shown in Table 1, we get $88$ apps from various categories such as “Health & Fitness”(iTunes App Store, 2024b) and “Developer Tools” (iTunes App Store, 2024a). These apps include $1414$ APIs, including all of $556$ APIs involved in $7627$ shortcuts. For more details about the extraction, please refer to the Appendix A.1.
@@ 3.2 Dataset Construction
(“a” in Figure 2) include the “API description” (“a.4”), “API name” (“a.1”), “parameter names” (“a.2”), “parameter types” (“a.1”), “default value” (“a.2”), “return value type” (“a.3”), and “return value name” (“a.3”). The field names in [...] in Figure 2 more details about \${filename}.actionsdata, \${filename}.intentdefinition, and WFActions.json, please refer to the Appendix A.2. In existing benchmarks / datasets (Table 1), the “parameter types” (“a.1”) and “return value types” (“a.3”) are composed of primitive data types such as int and string. In addition to primitive data types, APIs in ShortcutsBench also include “enum” or “advanced data types”. Enum is composed of “the class name” and “the possible value”, with each value equipping a “value name”. We also provide the agent with a description of the “enum” in the API information. Advanced data types, such as the model (“a.1”) in app chatGPT (OpenAI, 2024), include three String types named identifier, title, and subtitle. We can comprehend them through their “type name” and “type description”.
@@
Generating queries. As shown in Figure 1, existing works construct query and action sequences based on available APIs. In contrast, we construct queries based on existing action sequences and APIs. When constructing a query for a specific action sequence, we need to understand the functional description of the action sequence (“e” in Figure 2) and detailed information about the involved APIs (“a” in Figure 2). With this information, we can generate higher-quality queries. To ensure the quality of the generated queries, we also leverage the unique advantage of shortcuts: the natural language workflow descriptions (“b.7” / “c” in Figure 2). By inputting these intuitive natural language descriptions into an LLM, we can generate more accurate queries. When generating queries, we also require the model to naturally include primitive data type parameters and enum data type needed for API calls in generated queries. This helps us evaluate the agent’s ability to fill in primitive parameters in Section 3.3.2. To ensure the quality of generated queries, we use the state-of-the-art LLM, GPT-4o (OpenAI, 2024b), to generate the queries. The prompt templates we used to generate queries can be found in the Appendix A.2.
@@ 3.3 Task Definition and Metrics
Further Processing. Considering the context limitations of LLMs, we excluded shortcuts longer than $30$ and parts using the API is.workflow.actions.runworkflow to call other shortcuts. While these shortcuts remain in our open-source dataset, they will not be included in the subsequent evaluation. We aim to study the performance of agents on queries of varying difficulties. As shown in Table 2, we categorize ShortcutsBench into $4$ difficulty levels and $8$ task types based on $\left|aseq_{i}\right|$ and “shortcut type” (Section 3.1), respectively. For more details, please refer to the Appendix A.3. When calculating the length, for branching actions like is.workflow.actions.conditional, we consider the longest branch as the length. Additionally, we ignore the lengths of looping actions like is.workflow.actions.repeat.count and special actions such as is.workflow.actions.comment. Due to the presence of branching actions, the average number of APIs involved when $p=1$ is greater than one, specifically $1.17$. For a detailed process, please refer to the Appendix A.3. The number of shortcuts in each level is denoted as $n_{p}$. Each query and action sequence is referred to as $q_{p,i}$ and $aseq_{p,i}$, with $1\leq p\leq 4$ and $1\leq i\leq n_{p}$.
@@ 3.3.1 Performance about API Selection
Following existing work (Huang et al., 2023; Patil et al., 2023; Li et al., 2023b; Xu et al., 2023; Schick et al., 2024; Hao et al., 2024), we use the accuracy of API selection as the metric. The accuracy is calculated as the number of correct API selections $m_{p}$ divided by $n_{p}$. Specifically, each time we predict an action $b_{j},1\leq j\leq|aseq_{i}|$, we provide the agent with all the correct historical actions $\{a_{1},a_{2},...,a_{j-1}\}$. We then require the agent to predict the next action. All actions predicted by the agent form the prediction sequence $bseq_{p,i}$. This method is similar to the next token prediction (NTP) in LLMs, effectively preventing a cascade of errors in subsequent action predictions due to a single incorrect prediction. During the prediction, when encountering special actions such as branching and looping, we skip predicting these actions and directly add them to the historical actions. For more details, please refer to Appendix A.4. We chose API selection accuracy over the final result for the following two additional reasons:
@@ 3.3.2 Effectiveness of API Parameter Value Filling
  • Static Parameters Preset: These are static parameters that users provide as default inputs of the action. These static parameters typically include primitive data types such as String and Integer, as well as custom Enum defined by app developers. When the query explicitly specifies a parameter that can be used as a static parameter, we expect the agent to accurately fill in the parameter values according to the user’s query and the API’s definition. When generating queries, we have already required the LLM to naturally include primitive and enumerated data types (Section 3.2). To further ensure that the corresponding parameters are indeed included in the queries during evaluation, we used the LLM to filter these parameters further, ensuring their presence in the queries. Detailed prompts can be found in the Appendix A.5.
@@ 3.3.3 Recognition of Need for Input
In this section, we aim to investigate the ability of existing API-based agents to ask systems or users for necessary information to resolve the missing information. This missing information can come from the system like clipboard (Clipboard), input files (ExtensionInput), and the current date (CurrentDate) or from the user (Ask(Apple-Inc., 2024a). For example, a parameter named tags is usually represented in a shortcut as "tags":{"Value":{"Type": "Ask"}}, where "Type": "Ask" indicates that the parameter will prompt the user for input. For more details, please refer to Appendix A.6. We use the proportion of correctly identified parameters to evaluate the agent’s ability to recognize the need for input from the system or the user. Let $n_{s}$ be the number of queries, $aska_{i}$ be the number of times the need from the system or the user appears in $aseq_{i}$, $askt_{i}$ be the number of times the need from the system or the user appears in $bseq_{i}$. The accuracy of ask for necessary information can be calculated as $Acc_{afni}=askt_{i}/aska_{i}$.
@@ 4.1 Setup
Prompt Template. Following existing work (Huang et al., 2023; Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024; Hao et al., 2024), we slightly modified the ReACT (Yao et al., 2022) templates to construct the API-based agents. Specifically, we added prompts related to shortcuts, such as the types of fillable parameters and the meanings of special statements like branches. For all $3$ research questions (RQs), we use the same prompt templates. An agent should correctly select APIs, fill in parameters, and be aware of the need to request necessary information from the system or user at appropriate times. Please refer to Appendix A.7 for more details.
@@ 4.2 Result Analysis
  • Agents built using open-source LLMs now perform comparably to closed-source models on lower-difficulty tasks but still lag on higher-difficulty tasks. From Figure 4 we know that open-source LLMs $>=70$B match the performance of closed-source LLMs from the first $3$ difficulty tasks, significantly outperforming GPT-3.5-turbo. However, they still lag behind closed-source LLMs in handling complex tasks at the $4$-th level. Moreover, the price of open-source LLMs is significantly lower than that of GPT-3.5-turbo. For more details, please refer to Appendix A.8.